A 1.005 g sample of an unknown alkaline-earth metal was allowed to react with a volume of chlorine gas that contains 1.91 × 1022 Cl2 molecules. The resulting metal chloride was analyzed for chlorine by dissolving a 0.436 g sample in water and adding an excess of AgNO3(aq) to give a precipitate of 1.126 g of solid AgCl.

  1. What is the percent Cl in the alkaline-earth chloride?
  2. What is the identity of the alkaline-earth metal?
  3. Write balanced equations for all chemical reactions.
  4. In the reaction of the alkaline-earth metal with chlorine, which reactant is in excess and how many grams of the reactant remain unreacted?
Answer #1

#.(a): Moles of solid AgCl formed = mass of AgCl / molar mass of AgCl = 1.126 g / 143.32 g/mol = 0.007856545 mol AgCl

Since 1 mole AgCl contains 1 mol Cl, moles of Cl in 0.436 g metal chloride sample = 0.007856545 mol Cl

=> Grams of Cl present in 0.436 g metal chloride sample

= 0.007856545 mol Cl * Atomic mass of Cl = 0.007856545 mol*35.453 g/mol

= 0.27854 g Cl

=> Percent Cl in the alkaline-earth chloride = ( 0.27854 g / 0.436 g)*100 = 63.88 % (Answer)

#.(b):

Grams of metal in 0.436 g metal chloride sample = 0.436 g - 0.27854 g = 0.157462 g M

Since this is an alkaline earth metal, the chemical formula of the metal chloride will be MCl2.

i.e 1 mol metal atom contains 2 mol Cl atom.

Hence moles of metal in 0.436 g MCl2 =  0.007856545 mol Cl * (1 mol M / 2 mol Cl) = 0.00392827 mol M

=> Molar mass of metal = mass/mol = 0.157462 g M / 0.00392827 mol M = 40.08 g/mol

Since 40.08 g/mol is the atomic mass of Calcium, the identity of the alkaline earth metal is CaCl2 (Answer)

#.(c):

Balanced equations:

Ca(s) + Cl2(g) ---> CaCl2(s)

CaCl2(aq) + 2AgNO3(aq) -----> 2AgCl(s) + Ca(NO3)2(aq)

(d): Moles of Ca = 1.005 g / 40.08 g/mol = 0.02507485 mol

Moles of Cl = 1.91*1022 * (1 mol / 6.022*1023) = 0.031717 mol

Since the mole ratio is 1:1 and Ca has smaller mole, Ca is the limiting reactant.

and Cl2(g) is in excess.

Moles of Cl2(g) is in excess = 0.031717 mol - 0.02507485 mol = 0.006642 mol Cl2(g)

Grams remain unreacted = 0.006642 mol Cl2(g) * 70.9 g/mol = 0.471 g (Answer)